# Satellite Autonomous Clock Fault Monitoring with Inter-Satellite Ranges Using Euclidean Distance Matrices

### v2

@@ 5.1.1 Orbits and Links
    GPS constellation of 31 satellites: The satellite orbits are initialized using the Two-Line Element (TLE) data from 2025/4/3 from the CelesTrak website (celestrak_gp_data_formats_2025). The 3-dimensional plots of the constellation and the available ISLs at $t=0$ are shown in Figure 10(a).
@@
    A notational lunar constellation of 17 satellites: We considered a hybrid constellation, where 9 satellites form a 9/3/1 Walker-Delta constellation and 8 satellites are deployed in elliptical lunar frozen orbits (ELFO). The constellation is based on the design proposed by kang2025design, which provides a good coverage of the lunar south pole, while maintaining a good global coverage. The orbital elements of the constellation are shown in Table 5.1.1, and the 3-dimensional plots of the constellation and the available ISLs at $t=0$ are shown in Figure 10(c).
For both constellations, the orbits are propagated in the two-body propagator without considering perturbations. The availability of the links is calculated assuming that the links are available when it is not occulted by the central body, and the angle between the line of sight and the satellite-body center vector is less than $\phi_{max}$, as illustrated in Figure 9. The cutoff angle $\phi_{max}$ is set to 60 degrees for the Earth case (Zhang2024) and 80 degrees for the Moon case. To compute body occultation, we set an additional mask of 1000 km for the Earth case, and 100 km for the Moon case to avoid the link being corrupted by ionospheric delays or terrain blockage. We assume that we can obtain the range measurements from all available links at each time step. The scheduling of the links will be investigated in future work.
The number of 5 cliques in the constellation with the number of subgraphs containing PRN 1, 10, and 15 is shown in Figure 10(b) and Figure 10(d) for the Earth and Moon cases, respectively. For the lunar constellation, the ELFO satellites have time windows where they have no self-contained subgraphs, where the faults at the satellite cannot be detected using the EDM method.
@@
(a) Orbital planes with satellite position (black points) and links (orange line) at the initial time epoch. The constellation consists of 17 satellites, 9 of which are in the Walker constellation with 3 planes (shown in blue), and 8 satellites are in the 2 ELFO planes (shown in red).
@@
{tblr}
[…]
$a$ [km] $e$ [] $i$ [deg] $\Omega$ [deg] $\omega$ [deg] $M$ [deg]
1 1 - 3 6215.0 0.0 39.2 110 90 [0, 120, 240]
2 4 - 6 6215.0 0.0 39.2 230 90 [40, 160, 280]
3 7 - 9 6215.0 0.0 39.2 350 90 [80, 200, 320]
4 10 - 13 6215.0 0.7 55.0 0 90 [0, 90, 180, 270]
5 14 - 17 6215.0 0.7 55.0 180 90 [0, 90, 180, 270]
(c) Orbital planes with satellite position (black points) and links (orange line) at the initial time epoch. The constellation consists of 17 satellites, 9 of which are in the Walker constellation with 3 planes (shown in blue), and 8 satellites are in the 2 ELFO planes (shown in red).
(d) The scatter plot of the number of self-containing 5-cliques for each satellite in the constellation over full orbit. The number of 5-cliques that contain PRN 1, 5, and 15 are shown in blue, orange, and green, respectively. The satellites in ELFO (PRN 10 - 17) experience a time window near the periapsis where they have zero 5-cliques, where faults on that satellite cannot be detected.
Figure 11: A hybrid lunar constellation around the Moon with 17 satellites
@@ 5.1.2 Parameters
  • Test methods: ‘Sum of residuals’, ‘Data snooping’, and ‘EDM’
  • Fault magnitudes: $\bar{b}$ = 1, 2, 5, 10, 20 m
@@
  • False alarm rates for the fault detection algorithms: $\alpha=0.001,0.01,0.05$
For each of these cases, we run 5000 Monte-Carlo analyses with randomly sampled initial times (within one orbital period) and a set of fault satellites. For all cases, we set the standard deviation of the two-way noise measurement error to $\sigma_{m}=0.5$ m, based on the results of the Beidou constellation (Zhang2024), and used $\eta_{\alpha}=1.5$ for the threshold margin.
@@ 5.1.3 Evaluation Metrics
True Positive Rate or Recall($\text{True Positive Rate or Recall}\ (TPR)$ $=\frac{TP}{TP+FN}$ (60)
False Positive Rate($\text{False Positive Rate}\ (FPR)$ $=\frac{FP}{FP+TN}$ (61)
$P_{4}\ \text{metric}\ (P_{4})$ $=\frac{4\cdot TP\cdot TN}{4\cdot TP\cdot TN+(TP+TN)\cdot(FP+FN)}$ (62)
where $TP,FN,FP,TN$ is the total number of true positives (fault satellite classified as fault), false negatives (fault satellite classified as non-fault), false positives (non-fault satellite classified as fault), and true negatives (non-fault satellite classified as non-fault). A higher TPR is desired not to miss fault satellites, and a lower TPR is desired to prevent false alarms, detecting normal satellites as a fault. $P_{4}$ value is the harmonic mean of four key conditional probabilities (precision, recall, specificity, and negative predictive value), where $P_{4}=1$ requires all of these values to be 1 (Sitarz_2023). $P_{4}$ value could be used as a metric to evaluate the “balance” between these metrics.
@@ 5.2.1 Case 1: GPS Constellation
The fault detection results for the GPS constellation with different combinations of parameters are shown in Figure 12. When there are no fault satellites, FPRs of all methods are lower than the user-defined threshold $\alpha$ = 0.001, 0.01, 0.05. The sum of residuals method and the EDM method show the lowest FRP, due to their conservative threshold setting (assuming maximum correlation between the residuals for the sum of residuals method, and inflating the threshold for the EDM method).
The TPR and FPR of the sum of residuals method improve as the ephemeris accuracy improves, as shown in Figure 11(b) - 11(d). When the ephemeris accuracy is high ($\sigma_{m}=1$ m), the algorithm achieves nearly perfect fault detection when the fault magnitude is larger than 5 m, while with lower accuracy ($\sigma_{m}=3$ m), it will require a larger fault magnitude (10 m) to achieve the same TPR.
When the fault ratio is 1 (all ISRs containing fault satellites are corrupted), the data snooping method performs the best among the three methods, due to its optimality of the test. The EDM method also showed TPR over 0.7 for a fault magnitude of 2 m, marking a better P4 score than the sum of residuals method.
However, when the fault ratio is 0.2 (a smaller number of links are biased by the clock jump, such as due to link scheduling), the data snooping method showed worse TPR than the EDM method and the sum of residuals method with $\sigma_{m}=1$ m, since the data snooping test assumes all fault-satellite containing ISRs are biased. The data snooping method also showed worse FPR as the fault magnitude increases. For a fault ratio of 0.2, the EDM method showed the highest P4 score at fault magnitudes $\leq$ 10m, while the sum of residuals method showed the highest P4 score at fault magnitudes $\geq$ 10m.
(a) The results for the Earth case with no faults.
(b) The results for the Earth case with 1 fault satellite, a false alarm rate of 0.001, and a fault ratio of 1.
(c) The results for the Earth case with 1 fault satellite, a false alarm rate of 0.01, and a fault ratio of 1.
(d) The results for the Earth case with 1 fault satellite, a false alarm rate of 0.01, and a fault ratio of 0.2.
Figure 12: The fault detection results for the Earth case with 0 or 1 fault satellite. When there is no fault, the FPR for all methods went below the user-defined threshold $\alpha$ = 0.001, 0.01, 0.05. The TPR and FPR of the sum of residuals method improve as the ephemeris accuracy improves and the fault magnitude increases. When the fault ratio equals 1, the data snooping has better TPR and FPR at lower fault magnitudes due to the optimality of the test. However, when the fault ratio is 0.2 (a smaller number of links are biased by the clock jump), the TPR of the data snooping method becomes worse than the other methods, since the test assumes all fault-satellite containing ISRs are biased.
(a) The results for the lunar case with no faults.
(b) The results for the lunar case with 1 fault satellite, a false alarm rate of 0.001, and a fault ratio of 1.
(c) The results for the lunar case with 1 fault satellite, a false alarm rate of 0.01, and a fault ratio of 1.
(d) The results for the lunar case with 1 fault satellite, a false alarm rate of 0.01, and a fault ratio of 0.2.
Figure 13: The fault detection results for the lunar case with 0 or 1 fault satellite. When there is no fault, the FPR of all methods went below the user-defined threshold $\alpha$ = 0.001, 0.01, 0.05. The TPR of the EDM and the data snooping methods caps around 0.91, due to the small number of available links near the perilune for the ELFO satellites making the test statistic less sensitive to the bias (the bias becoming less observable in the linear system). When the fault ratio is 0.2 (a smaller number of links are biased by the clock jump), the EDM method shows the highest P4 score at fault magnitudes $\leq$ 10m, while the sum of residuals method shows the highest P4 score at fault magnitudes $\geq$ 20m.
@@ 5.2.2 Case 2: Lunar Constellation
The metrics for the lunar cases are shown in Figure 13. Similar to the Earth GPS case, when there is no fault, the FPR of all methods are lower than the user-defined threshold $\alpha$ = 0.001, 0.01, 0.05, and the TPR and FPR of the sum of residuals method improve as the ephemeris accuracy improves and the fault magnitude increases.
When the fault ratio is 1, the TPR of the EDM and the data snooping methods caps around 0.91, due to the small number of available links near the perilune for the ELFO satellites, making the test statistic less sensitive to the bias. When the fault magnitudes are larger ($\geq$ 10 m), the sum of residuals method shows larger TPR and lower FPR than the other two methods, since the method directly compares the observed range to the expected range, and thus the performance does not depend on the overall geometry of the links (i.e., or the observability of the bias in the system). When the fault ratio is 0.2 (a smaller number of links are biased by the clock jump), the EDM method shows the highest P4 score at fault magnitudes $\leq$ 10m, while the sum of residuals method shows the highest P4 score at fault magnitudes $\geq$ 20 m. Overall, the EDM method shows the most balanced performance (P4 score) over different fault magnitudes and fault ratios for the lunar case.
@@ 6 Conclusion
In this paper, we proposed a new clock phase jump detection algorithm for satellite constellations using ISR measurements. The method can operate autonomously within the constellation. In addition, since the fault detection directly processes the measurement ISR measurements, it does not require prior knowledge of the satellite positions or clock biases, offering a flexible solution for LunaNet constellations, which may include nodes that are not dedicated for PNT.
In this paper, we first formulated the equations of the dual one-way ISR measurements, and modeled how they are affected by the clock phase jumps. We then introduced two baseline methods to detect the clock phase jumps, based on the residuals of the ISR measurements, which are the sum of residuals method and the data snooping method.
Then, we proposed a new fault detection algorithm based on the EDMs and GCEDMs derived from the ISR measurements. We proved several properties of the rank of the EDM and GCEDM, as well as the distribution of the 4th singular value of the GCEDM. Based on these properties, we demonstrated that satellite fault can be detected by analyzing the 4th singular values of the GCEMs derived from ISR measurements. The proposed fault detection algorithm computes the sum of the GCEDM’s 4th singular value for 5-clique subgraphs, and compares it to a threshold based on the chi-squared distribution of the 4th singular value under the presence of noise.
Through simulations of a GPS constellation and a notional lunar constellation, we compared the performance of the proposed method with the two baseline methods based on range residuals: the “sum of residuals” method and the “data snooping” method. The sum of residuals method performed well for larger fault magnitudes, but its detection power (TPR) was lower than the other two methods for smaller fault magnitudes. The data snooping method performed best among all methods when all ISRs containing fault satellites were corrupted, but its performance degraded when the biases were less observable (lunar scenario), and only a few ISRs were corrupted (e.g., due to link scheduling). The proposed EDM-based approach demonstrated balanced performance across different fault magnitudes and fault ratios for both GPS and lunar constellation scenarios.
This research marks a step towards establishing reliable PNT services for LunaNet. Future work will focus on link scheduling to optimize the number of ISRs available for fault detection under a constrained number of links, selecting effective subgraphs to reduce computation, looking into distributed implementation, and combining the EDM-based fault detection with residual-based fault detection methods to further enhance the robustness of the algorithm.
@@ Appendix A Computing the Test Threshold Using Imhof’s Method
In this section, we describe the procedure to obtain the critical value for the hypothesis test in section 3.1. The correlation between the two residuals that share the same satellites can be described as (xu2011gnss)
@@
Using the Imhof’s method, the cumulative distribution function (CDF) value of the test statistic $T_{i}=\sum_{j=1}^{l_{i}}\tilde{g}_{ij}^{2}$ is given as (imhof1961computing)
@@
where $\lambda_{k}(k=1,\ldots,l_{i})$ are the eigenvalues of the covariance matrix $\Sigma$. However, it is computationally expensive to evaluate equation (66) at every satellite at every timestep. Therefore, xu2011gnss suggests to use the conservative bound for the cross-terms of $\Sigma_{i,j}$, as follows
@@
$G\ \text{contains}\ m\ (1\leq m\leq k-1)\ \text{faulty nodes}$ (73)
Without the loss of generality, we can assume the indices of the faulty nodes are $1,\ldots,m$ for $m\leq k-1$. Let $G^{\prime}=G\ \backslash\ \{1,\ldots,m\}$ a graph without the $m\leq k-1$ faulty nodes. Since $G$ is $k$ vertex redundantly rigid, $G^{\prime}$ is rigid in $\mathbb{R}^{d}$. In addition, since $G^{\prime}$ contains no fault nodes, we have
@@
Geometrically, this is the common intersection of $\|\mathcal{N}(i)\|$ spheres in $\mathbb{R}^{3}$. It is known that every vertex in a $k$-vertex redundantly rigid graph must have degree at least $d+k-1$ (Jordan2022extremal). Since $m\leq k-1$, each fault node has at maximum $k-2$ edges with other fault nodes (which are not included in $G^{\prime}$), and therefore we have $\|\mathcal{N}(i)\|\geq(d+k-1)-(k-2)=3-1+2=4$.
@@
Since the Jacobian of $F$ can have rank at most 3, its image $F(\mathbb{R}^{3})\subset\mathbb{R}^{4}$ is at most a 3-dimensional immersed submanifold of $\mathbb{R}^{4}$. Any smooth embedding of a $k$-dimensional manifold into $\mathbb{R}^{n}$ has $n$-dimensional Lebesgue measure 0 whenever $k<n$ (GuilleminPollack1974), and since the law of random radii vector $\mathbf{r}$ is assumed continuous with respect to $4$-dimensional Lebesgue measure, the probability that equation (80) is satisfied is 0. This contradicts the statement that the graph $G$ is $d$-embeddable with the observed ranges, at probability 1. Since we derived a falsehood, the assumption (73) that the graph contains $11\leq m\leq k-$ 1 fault nodes is false, at probability 1. Therefore, if the graph is $k$-vertex rigid and is $d$-embeddable, and the graph contains at most $k-1$ faulty nodes, the graph does not contain any fault nodes, at probability 1. ∎
@@
As discussed in dokmanic2015euclidean, the noiseless- and faultess-EDM, $\mathbf{D}^{n,d,0}_{ij}=\|\mathbf{x}_{i}-\mathbf{x}_{j}\|^{2}$, based on a collection of points $\mathbf{X}\in\mathbb{R}^{d\times n}$ can be constructed with Equation (81).
@@
where $\mathrm{diag}(\mathbf{X}^{\top}\mathbf{X})\in\mathbb{R}^{n}$ is the vector formed from the diagonal entries of $\mathbf{X}^{\top}\mathbf{X}$. The first and last matrices of Equation (81) are rank 1 matrices by construction. These rank 1 matrix each contribute to the matrix rank except for degenerate cases where the points are equally spread apart ($\mathbf{X}\mathbf{1}=\mathbf{0}$). The middle term includes the Gram matrix $\mathbf{X}^{\top}\mathbf{X}$, which is rank $d$ except for degenerate cases where the points lie on a lower dimensional hyperplane (i.e., $\mathrm{rank}(\mathbf{X}^{\top}\mathbf{X})=2$ if the points lie on a plane). So, in line with dokmanic2015euclidean, $\mathrm{rank}(\mathbf{D}^{n,d,0})\leq d+2$, where the condition holds with equality outside of the aforementioned degenerate cases.

### v3

@@ 5.1.1 Orbits and Links
    GPS constellation of 31 satellites: The satellite orbits are initialized using the Two-Line Element (TLE) data from 2025/4/3 from the CelesTrak website [17]. The 3-dimensional plots of the constellation and the available ISLs at $t=0$ are shown in Figure 11(a).
@@
    A notational lunar constellation of 9 satellites: We consider a constellation of 9 satellites in 3 orbital planes of elliptical lunar frozen orbit (ELFO) with a 30-hour orbital period, as a potential extension of the Lunar Communications Relay and Navigation Systems (LCRNS) [26]. The orbital elements of the constellation are shown in Table 1, and the 3-dimensional plots of the constellation and the available ISLs at $t=0$ are shown in Figure 12(a).
For both constellations, the orbits are propagated in the two-body propagator without considering perturbations. The availability of the links is calculated assuming that the links are available when it is not occulted by the central body, and the angle between the line of sight and the satellite-body center vector is less than $\phi_{max}$, as illustrated in Figure 10. The cutoff angle $\phi_{max}$ is set to 60 degrees for the Earth case [33] and 90 degrees for the Moon case. To compute body occultation, we set an additional mask of 1000 km for the Earth case, and 100 km for the Moon case to avoid the link being corrupted by ionospheric delays or terrain blockage, respectively. We assume that we can obtain the range measurements from all available links at each time step. The scheduling of the links will be investigated in future work.
The scatter plot of the number of 5 cliques in the constellation containing each satellite is shown in Figure 11(b) and Figure 12(c) for the Earth and Moon cases, respectively. For the lunar constellation, we observe that the number of available 5-cliques is significantly lower (below 30), limiting the use of EDM-based fault detection from ISR alone. Therefore, for the lunar case, we used the ephemeris augmentation described in Section 4.3.3 to increase the number of subgraphs that can be used for fault detection. The number of fault-detectable 5-node subgraphs are shown in 12(c)
@@
(a) Orbital planes with satellite position (black points) and links (orange line) at the initial time epoch. The constellation consists of 31 satellites in six orbital planes.
@@
Plane PRN Semi-Major Axis EccentricityEccentricity Inclination RAAN Argument of Periapsis Mean Anomaly
$a$ [km] e[][] $i$ [deg] $\Omega$ [deg] $\omega$ [deg] $M$ [deg]
1 1 - 3 11314.7 $0.56$ 56.8 206.6 90 [0, 120, 240]
2 4 - 6 11314.7 $0.56$ 46.9 321.2 98.1 [40, 160, 280]
3 7 - 9 11314.7 $0.56$ 46.9 91.9 81.9 [80, 200, 320]
(a) Orbital planes with satellite position (black points) and links (orange line) at the initial time epoch. The constellation consists of 9 satellites in the Walker constellation with 3 planes
(b) The scatter plot of the number of self-containing 5-cliques for each satellite in the constellation. The number of 5-cliques that contain PRN 1, 2, and 3 are shown in blue, orange, and green, respectively.
(c) The scatter plot of the number of self-containing fault-detectable 5-sat subgraphs for each satellite in the constellation. All satellites have the same number of self-containing fault-detectable subgraphs.
Figure 12: A lunar constellation around the Moon with 9 satellites
@@ 5.1.2 Parameters
  • Test methods: “ephemeris comparison”, ‘Data snooping’, and ‘EDM’
  • Fault magnitudes: $\bar{b}$ = 2, 4, …, 20 m
@@
  • Ephemeris accuracy: $\sigma_{m}=\{$1\text{\,}\mathrm{m}$,$2\text{\,}\mathrm{m}$\$\} for Earth, $\sigma_{m}=\{$2\text{\,}\mathrm{m}$,$3\text{\,}\mathrm{m}$,$4\text{\,}\mathrm{m}$\$\} for Moon.
  • Input probability of false alarm for the fault detection algorithms:
    $\alpha=0.001,0.002,0.003,0.005,0.008,0.013,0.022,0.036,0.06,0.1$
For each of these cases, we run 5000 Monte-Carlo analyses with randomly sampled initial times (within one orbital period) and a set of fault satellites. For all cases, we set the standard deviation of the two-way noise measurement error to $\sigma_{m}=0.5$ m, based on the results of the Beidou constellation [33]. The inflation rate of $\eta_{\alpha}=1.5$ is used for the Earth scenario and $\eta_{\alpha}=5.0$ is used for the lunar scenario.
@@ 5.1.3 Evaluation Metrics
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\text{Probability of Missed Detection}\ P_{md}}$ $=\frac{FN}{TP+FN}$ (62)
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\text{Probability of False Alarm}\ P_{fa}}$ $=\frac{FP}{FP+TN}$ (63)
where $TP,FN,FP,TN$ are the total number of true positives (fault satellite classified as fault), false negatives (fault satellite classified as non-fault), false positives (non-fault satellite classified as fault), and true negatives (non-fault satellite classified as non-fault). A lower $P_{md}$ is desired not to miss fault satellites, and a lower $P_{fa}$ is desired to prevent false alarms, detecting normal satellites as a fault.
@@ 5.2.1 Case 1: GPS Constellation
Figure 13 summarizes the fault detection performance for the GPS constellation under different parameter combinations.
a) No fault
When no satellites are faulty, the false alarm probability $P_{fa}$ of all methods remains below the user-specified threshold $\alpha=0.001$$0.1$. Among the three approaches, the data snooping method and the EDM-based method achieve the lowest $P_{fa}$, indicating strong robustness in fault-free conditions.
b) Single faulty satellite with fault ratio = 1
When the fault ratio is 1, meaning that all inter-satellite ranges (ISRs) involving the faulty satellite are corrupted, the data snooping method yields the best performance. Owing to the optimality of its statistical test under this assumption, it achieves nearly perfect fault detection ($P_{fa}=P_{md}=0$) for fault magnitudes as small as 2 m, as shown in Figure 13(e). The $P_{fa}$$P_{md}$ trade-off curve of the EDM test lies between those of data snooping and the ephemeris comparison method, as illustrated in the same figure.
The performance of the ephemeris comparison method improves as the ephemeris accuracy increases, as shown in Figure 13(b). For $\alpha=0.001$ and ephemeris accuracy $\sigma_{m}=1$ m, perfect fault detection ($P_{fa}=P_{md}=0$) is achieved when the fault magnitude exceeds 5 m. In contrast, with lower ephemeris accuracy ($\sigma_{m}=2$ m), a larger fault magnitude (approximately 8 m) is required to drive both $P_{fa}$ and $P_{md}$ to zero.
c) Single faulty satellite with fault ratio = 0.2
When the fault ratio is reduced to 0.2, corresponding to a scenario in which only a subset of links is biased (e.g., due to link scheduling), the assumptions underlying the data snooping test are violated. Consequently, data snooping is no longer optimal. Nevertheless, for small fault magnitudes, it still provides the most favorable $P_{fa}$$P_{md}$ trade-off among the three methods.
For larger fault magnitudes (e.g., $\geq 15$ m), the ephemeris comparison method outperforms the other two approaches. This behavior arises because, when corrupted edges are sparse, both the EDM and data snooping methods exhibit a higher probability of misattributing the fault to non-faulty satellites. As in the previous cases, the EDM method consistently demonstrates intermediate performance between ephemeris comparison and data snooping.
(a) [Earth] Probability of false alarm when there are no faults.
(b) [Earth] 1 fault, $\alpha$: 0.001, fault ratio: 1
(c) [Earth] 1 fault, $\alpha$: 0.001, fault ratio: 0.2
(d) Earth] 1 fault, $\alpha$: 0.013, fault ratio: 0.2
(e) [Earth] $P_{fa}$ vs $P_{md}$ for different fault magnitudes (from left to right: 2m, 4m, 8m, 16m) and input $P_{fa}$ (varied inside each subplot)
Figure 13: The fault detection results for the Earth case with 0 or 1 fault satellite. When there is no fault, the $P_{fa}$ of all methods went below the user-defined $P_{fa}$ $\alpha$ = 0.001 - 0.01. When the fault ratio is 1.0, the data snooping showed the lowest $P_{md}$ and $P_{fa}$, followed by the EDM approach and ephemeris approach.
(a) [Moon] Probability of false alarm when there are no faults.
(b) [Moon] 1 fault, $\alpha$: 0.001, fault ratio: 1
(c) [Moon] 1 fault, $\alpha$: 0.001, fault ratio: 0.2
(d) [Moon] 1 fault, $\alpha$: 0.013, fault ratio: 0.2
(e) [Moon] $P_{fa}$ vs $P_{md}$ for different fault magnitudes (from left to right: 2m, 4m, 8m, 16m) and input $P_{fa}$ (varied inside each subplot)
Figure 14: The fault detection results for the lunar case with 0 or 1 fault satellite. When there is no fault, the $P_{fa}$ of all methods went below the user-defined $P_{fa}$ ($\alpha$ = 0.001 - 0.1). For fault ratio of 1.0, the data snooping showed the lowest $P_{md}$ and $P_{fa}$, followed by the EDM approach and then the ephemeris approach. When the fault ratio is 0.2, the EDM and Ephemeris approach provides better $P_{fa}$ and $P_{md}$ for large fault magnitudes.
@@ 5.2.2 Case 2: Lunar Constellation
Figure 14 summarizes the fault detection performance for the lunar constellation under different parameter combinations. Since the ephemeris augmentation strategy described in Section 4.3.3 is employed for the EDM approach, its detection performance also becomes explicitly dependent on the assumed ephemeris accuracy.
a) No fault
Consistent with the GPS case, when no satellites are faulty, the false alarm probability $P_{fa}$ of all methods remains below the user-specified threshold $\alpha=0.001$$0.1$. Among the three approaches, the data snooping method and the EDM-based method achieve the lowest $P_{fa}$, indicating robust behavior under fault-free lunar operating conditions.
b) Single faulty satellite with fault ratio = 1
When the fault ratio is 1.0, the data snooping method again yields the best overall performance, owing to the optimality of its statistical test under the assumption that all links associated with the faulty satellite are corrupted. However, due to the reduced number and sparsity of inter-satellite links in the lunar constellation compared to the GPS case, a larger fault magnitude (approximately $\geq 10$ m) is required for data snooping to achieve nearly perfect fault detection ($P_{fa}\approx P_{md}\approx 0$), as shown in Figures 14(b) and 14(e).
As in the terrestrial case, the $P_{fa}$$P_{md}$ trade-off curve of the EDM approach lies between those of data snooping and the ephemeris comparison method. The performance of the ephemeris comparison method improves with increasing ephemeris accuracy, although its sensitivity to ephemeris errors is more pronounced in the lunar case due to weaker geometric redundancy.
c) Single faulty satellite with fault ratio = 0.2
When the fault ratio is reduced to 0.2, corresponding to scenarios in which only a subset of inter-satellite links is corrupted, the assumptions underlying the data snooping test are violated. As a result, the performance of data snooping degrades significantly, with elevated $P_{md}$ observed even for moderate fault magnitudes.
In contrast, both the EDM and ephemeris comparison methods exhibit improved robustness in this sparse-fault regime. For larger fault magnitudes (e.g., $\geq 8$–10 m), these two approaches achieve lower $P_{md}$ while maintaining $P_{fa}$ near or below the prescribed threshold. The EDM method consistently demonstrates intermediate performance, outperforming data snooping for sparse faults while remaining less sensitive to ephemeris inaccuracies than the ephemeris comparison method when high-accuracy ephemerides are not available.
@@ 6 Conclusion
This paper presented a novel framework for detecting satellite clock phase jumps using inter-satellite range measurements, grounded in rigid graph theory and Euclidean distance matrix analysis. The proposed method models the constellation as a weighted graph and evaluates the realizability of vertex-redundantly rigid subgraphs to identify clock jumps that manifest as structured biases in the inter-satellite ranges.
We established the necessary graph-topological conditions for fault detectability and derived analytical expressions for the distribution of the singular values of the geometric-centered Euclidean distance matrix (GCEDM) under both nominal and faulty conditions. These results enabled the development of statistically rigorous hypothesis tests with controllable false-alarm probabilities and analytically characterized detection performance. To address sparse measurement geometries, we introduced an ephemeris-augmented formulation that increases fault detectability when inter-satellite connectivity is limited, at the cost of increased sensitivity to ephemeris uncertainty. Simulation studies using both a terrestrial GPS constellation and a notional lunar constellation demonstrated that the proposed method remains effective across a wide range of fault magnitudes, link topologies, and operational scenarios.
Overall, the proposed approach offers a geometry-driven perspective on range-based fault detection through the lens of graph rigidity, complementing conventional residual-based methods. Future work will investigate extensions that incorporate modifications to data-snooping methods, multi-fault detection, and integration with link scheduling to maximize fault observability and detection performance.
@@ Appendix A Computing the Test Threshold Using Imhof’s Method
In this section, we describe the procedure to obtain the critical value for the hypothesis test in section 3.1. The correlation between the two residuals that share the same satellites can be described as [32]
@@
Using the Imhof’s method, the cumulative distribution function (CDF) value of the test statistic $T_{i}=\sum_{j=1}^{l_{i}}\tilde{g}_{ij}^{2}$ is given as [12]
@@
where $\lambda_{k}(k=1,\ldots,l_{i})$ are the eigenvalues of the covariance matrix $\Sigma$. However, it is computationally expensive to evaluate equation (67) at every satellite at every timestep. Therefore, [32] suggests to use the conservative bound for the cross-terms of $\Sigma_{i,j}$, as follows
@@
[…]
Without the loss of generality, we can assume the indices of the faulty nodes are $1,\ldots,m$ for $m\leq k-1$. Let $G\{$$^{\prime}=G\ \backslash\ \{1,\ldots,m\}$ a graph without the $m\leq k-1$ faulty nodes. Since $G$ is $k$ vertex redundantly rigid, $G^{\prime}$ is rigid in $\mathbb{R}^{d}$. In addition, since $G^{\prime}$ contains no fault nodes, we have
@@
Geometrically, this is the common intersection of $\|\mathcal{N}(i)\|$ spheres in $\mathbb{R}^{3}$. It is known that every vertex in a $k$-vertex redundantly rigid graph must have degree at least d+k1d+k-1 [16]. Since $m\leq k-1$, each fault node has at maximum k2k-2 edges with other fault nodes (which are not included in $G^{\prime}$), and therefore we have $\|\mathcal{N}(i)\|\geq(d+k-1)-(k-2)=3-1+2=4$.
@@
Since the Jacobian of $F$ can have rank at most 3, its image $F(\mathbb{R}^{3})\subset\mathbb{R}^{4}$ is at most a 3-dimensional immersed submanifold of $\mathbb{R}^{4}$. Any smooth embedding of a $k$-dimensional manifold into $\mathbb{R}^{n}$ has $n$-dimensional Lebesgue measure 0 whenever $k<n$ [8], and since the law of random radii vector $\mathbf{r}$ is assumed continuous with respect to $4$-dimensional Lebesgue measure, the probability that equation (81) is satisfied is 0. This contradicts the statement that the graph $G$ is $d$-embeddable with the observed ranges, at probability 1. Since we derived a falsehood, the assumption (74) that the graph contains $1\leq m\leq k-1$ fault nodes is false, at probability 1. Therefore, if the graph is $k$-vertex rigid and is $d$-embeddable, and the graph contains at most k1k-1 faulty nodes, the graph does not contain any fault nodes, at probability 1. ∎
@@
As discussed in [7], the noiseless- and faultess-EDM, $\mathbf{D}^{n,d,0}_{ij}=\|\mathbf{x}_{i}-\mathbf{x}_{j}\|^{2}$, based on a collection of points $\mathbf{X}\in\mathbb{R}^{d\times n}$ can be constructed with Equation (82).
@@
where $\mathrm{diag}(\mathbf{X}^{\top}\mathbf{X})\in\mathbb{R}^{n}$ is the vector formed from the diagonal entries of $\mathbf{X}^{\top}\mathbf{X}$. The first and last matrices of Equation (82) are rank 1 matrices by construction. These rank 1 matrix each contribute to the matrix rank except for degenerate cases where the points are equally spread apart ($\mathbf{X}\mathbf{1}=\mathbf{0}$). The middle term includes the Gram matrix $\mathbf{X}^{\top}\mathbf{X}$, which is rank $d$ except for degenerate cases where the points lie on a lower dimensional hyperplane (i.e., $\mathrm{rank}(\mathbf{X}^{\top}\mathbf{X})=2$ if the points lie on a plane). So, in line with [7], $\mathrm{rank}(\mathbf{D}^{n,d,0})\leq d+2$, where the condition holds with equality outside of the aforementioned degenerate cases.